# VIR-Bench: Evaluating Geospatial and Temporal Understanding of MLLMs via Travel Video Itinerary Reconstruction

### v1

@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) Liu et al. (2023); Li et al. (2024a); Lin et al. (2024); OpenAI (2024a) have improved remarkable capabilities in video understanding. Lately, research attention has shifted toward evaluating the spatial and temporal reasoning abilities of MLLMs, prompting the proposal of new benchmarks Grauman et al. (2022); Chandrasegaran et al. (2024); Jia et al. (2024); Yang et al. (2025); Lin et al. (2025). However, existing benchmarks primarily focus on micro-scale scenarios, such as indoor scenes or short-range outdoor activities, leaving macro-scale geospatial scenarios, namely, long-distance travel activities involving multi-day footage across multiple cities, largely unexplored. We argue that long-horizon geospatial-temporal reasoning is essential for next-generation MLLMs, as numerous real-world applications, such as embodied AI planning, navigation, and autonomous driving, heavily rely on these capabilities.
To address this gap, we introduce VIR-Bench, a benchmark to evaluate long-range geospatial-temporal understanding via itinerary reconstruction from travel vlog videos. The core output is a directed visiting order graph Yamamoto et al. (2025): nodes represent locations at three granularities (prefecture, city, and point of interest (POI)) and edges represent two relations, inclusion for spatial hierarchy and transition for temporal adjacency. By decomposing the task into two sub-tasks: (1) node prediction, identifying all locations visited; and (2) edge prediction, inferring geographic inclusion relations and temporal transition relations among visited locations, VIR-Bench enables separate evaluation of geospatial and temporal intelligence. Because the footage is mostly egocentric or selfie-style, models must construct a holistic understanding from partial views, which further stresses geospatial–temporal reasoning. The dataset comprises 200 travel vlogs filmed across Japan, a major inbound tourism destination, each accompanied by a manually annotated and double-reviewed visiting order graph.
@@ 2.1 Video Benchmarks
With the rapid advancement of MLLMs, recent video understanding benchmarks have increasingly prioritized evaluating models’ spatial and temporal reasoning capabilities. For instance, Ego4D Grauman et al. (2022) facilitates the evaluation of models’ comprehension of past and future events by utilizing curated temporal data, while HourVideo Chandrasegaran et al. (2024) examines the performance of models in understanding extended-duration video content. VSI-Bench Yang et al. (2025) assesses a model’s ability to infer 3D scene layouts from 2D video inputs, while OST-Bench Lin et al. (2025) evaluates spatial-temporal understanding by requiring models to explore and interpret information within a 3D space. CityGuessr Kulkarni et al. (2024) introduces a video-based benchmark for assessing geo-localization using driving videos, while UrbanVideo-Bench Zhao et al. (2025) targets the embodied cognitive abilities of MLLMs within urban 3D environments using drone-collected footage. Nevertheless, most existing benchmarks primarily feature indoor scenarios or short-distance outdoor movements, lacking extensive long-distance traversal, such as inter-city journeys. Consequently, these benchmarks are insufficient for thoroughly evaluating the geospatial-temporal intelligence of MLLMs. In contrast, VIR-Bench specifically addresses this gap by comprehensively assessing video understanding capabilities across extended spatial (e.g., from Tokyo to Osaka, Hokkaido to Kyoto) and temporal (spanning multiple days) scales.
@@ 2.2 Itinerary Extraction
Researchers in natural language processing have substantially studied the task of extracting travel trajectories from text Drymonas and Pfoser (2010); Kaushik et al. (2017); Haris and Gan (2021); Yamamoto et al. (2025). A representative study by Yamamoto et al. (2025) introduces a visiting order graph designed to capture relationships among visited locations and provides a benchmark dataset for training and evaluating itinerary extraction models. In multimodal settings, Pang et al. (2011) proposes a framework aimed at summarizing travelogues by integrating text and images from blogs. More recently, Rosa (2024) leverages MLLMs to perform structured entity extraction from travel videos, while Zhuang et al. (2024) tackles the inverse problem by generating vlogs with diverse travel scenes. This study advances this line of work by providing, to our knowledge, the first systematic investigation into extracting and reconstructing itineraries directly from videos, establishing a new benchmark for video-centric geospatial and temporal understanding.
@@ 2.3 Itinerary Generation
The complexity of manual trip planning has driven research into automated itinerary generation. Initial approaches were often based on optimization problems like the Tourist Trip Design Problem and classic machine learning Gavalas et al. (2014); Chen et al. (2016); He et al. (2019); Carrillo et al. (2023). More recently, Large Language Models (LLMs) have enabled more sophisticated and flexible frameworks Chen et al. (2024); Xie et al. (2024). Current research trends include developing novel reasoning paradigms Gui et al. (2025), creating hybrid systems that combine LLMs with classical planners de la Rosa et al. (2024), and enhancing personalization through user model integration and interactive feedback Singh et al. (2024); Chen et al. (2024); Otaki and Baba (2025).
To rigorously evaluate these methods, various benchmarks have been developed. Notable studies include real-world planning Xie et al. (2024), fine-grained spatio-temporal planning Chaudhuri et al. (2025), and assessing personalization Singh et al. (2024).
@@ 3.1 Visiting Order Graph
We adopt and refine the definition of a visiting order graph introduced by Yamamoto et al. (2025). A simplified example is shown in Figure 2.
@@
To prevent cycles in the graph, we treat multiple visits to the same location as distinct nodes. Following Yamamoto et al. (2025), we also introduce a special “Overlap” edge to handle POIs that are geographically overlapping but cannot be represented through inclusion edges.
@@ 3.2 Data Annotation
After annotation, we retrieved detailed information of each POI including name, address and categories using Google Places API. Using the timestamped POIs, we then constructed a visiting order graph for each video (e.g., Figure 2). We also conducted a quality check of the generated graphs and corrected errors by rerunning the retrieval step manually when POIs were incorrectly annotated or matched. This pipeline yielded VIR-Bench, a dataset of 200 travel videos paired with their corresponding visiting order graphs. In total, 3,689 POIs were identified across 43 of Japan’s 47 prefectures. Detailed annotation guidelines and dataset statistics are provided in Appendix A.
@@ 4.2 Benchmark Models
We evaluate the performance of mainstream MLLMs on VIR-Bench, including both open-weight models (VideoLLaMA3 Zhang et al. (2025), LLaVA-Video Zhang et al. (2024)), InternVL3 Zhu et al. (2025), Qwen2.5-VL Bai et al. (2025)) and proprietary models (GPT-4.1 OpenAI (2025a), o4-mini OpenAI (2025b), Gemini-2.5-Flash and Pro Comanici et al. (2025a)). All models are evaluated in a zero-shot setting. We use as many input frames as permitted by each model’s interface or pre-training setup; only the Gemini models accept audio input, and full details appear in Appendix B.1.
@@ 4.4 Main Results
When comparing models of different sizes within the same families (LLaVA-Video, InternVL3, Qwen2.5-VL), we observe steady, scale-driven gains on node prediction, whereas edge prediction shows a sharp jump; for example, transition F1 improves by about $16\times$ from Qwen2.5-VL-7B to Qwen2.5-VL-72B. This pattern reflects the task demands: node prediction is a localized, single-point task that primarily relies on geospatial knowledge encoded in the models, whereas edge prediction requires a holistic view of the itinerary and thus benefits more from larger models with stronger long-context and temporal reasoning. An exception is LLaVA-Video-72B, which shows minimal improvement over LLaVA-Video-7B in POI node prediction. This is likely due to the limited geographic coverage in the LLaVA-OneVision training data Li et al. (2024b). In contrast, models like Qwen2.5-VL demonstrate strong geo-localization capabilities; in our internal evaluations, Qwen2.5-VL was able to accurately predict POI coordinates, indicating extensive pretraining on geographic data.
@@ 5.1 Task Definition
The agent system takes a list of POIs, a video and planning constraints as input. While this list of POIs could be the output of the node prediction task, we use the POI list from the video annotation in this experiment to isolate the evaluation from model performance. The video provides richer information for the planning process. The constraints consist of the number of companions, travel duration and travel budget inspired by previous work Xie et al. (2024).
@@ 5.3.2 Evaluation Setup
We also evaluate the system’s POIs selection. Using GPT-4o OpenAI (2024b), we extract the POIs mentioned in the generated plans and treat them as selected. We then compare selected versus unselected POIs in terms of their on-screen duration in the video and their Google Maps ratings.
@@ A.1.3 Trajectory Annotation
We use Vidat Zhang et al. (2020), a browser-based video annotation tool. Before starting, we strongly recommend watching the official 7-minute tutorial video.
@@ B.1 Model Cards
VideoLLaMA3-7B 180 - - Zhang et al. (2025) DAMO-NLP-SG/VideoLLaMA3-7B
LLaVA-Video-7B 64 - - Zhang et al. (2024) lmms-lab/LLaVA-Video-7B-Qwen2
@@
InternVL3-8B 64 - - Zhu et al. (2025) OpenGVLab/InternVL3-8B
@@
Qwen2.5-VL-7B 256 - - Bai et al. (2025) Qwen/Qwen2.5-VL-7B-Instruct
@@
GPT-4.1 256 - - OpenAI (2025a) gpt-4.1-2025-04-14
o4-mini 256 32768 - OpenAI (2025b) o4-mini-2025-04-16
Gemini-2.5-Flash 1 fps - 1 kbps Comanici et al. (2025b) gemini-2.5-flash

### v2

@@ 1 Introduction
Recent advances in multimodal large language models (MLLMs) liu2023visualinstructiontuning; li2024llavanextinterleavetacklingmultiimagevideo; Lin_2024_CVPR; openai2024gpt4technicalreport have improved remarkable capabilities in video understanding. Lately, research attention has shifted toward evaluating the spatial and temporal reasoning abilities of MLLMs, prompting the proposal of new benchmarks grauman2022ego4dworld3000hours; chandrasegaran2024hourvideo; jia2024sceneversescaling3dvisionlanguage; yang2025thinking; lin2025ost. However, existing benchmarks primarily focus on micro-scale scenarios, such as indoor scenes or short-range outdoor activities, leaving macro-scale geospatial scenarios, namely, long-distance travel activities involving multi-day footage across multiple cities, largely unexplored. We argue that long-horizon geospatial-temporal reasoning is essential for next-generation MLLMs, as numerous real-world applications, such as embodied AI planning, navigation, and autonomous driving, heavily rely on these capabilities.
To address this gap, we introduce VIR-Bench, a benchmark to evaluate long-range geospatial-temporal understanding via itinerary reconstruction from travel vlog videos. The core output is a directed visiting order graph yamamoto-etal-2025-graph: nodes represent locations at three granularities (prefecture, city, and point of interest (POI)) and edges represent two relations, inclusion for spatial hierarchy and transition for temporal adjacency. By decomposing the task into two sub-tasks: (1) node prediction, identifying all locations visited; and (2) edge prediction, inferring geographic inclusion relations and temporal transition relations among visited locations, VIR-Bench enables separate evaluation of geospatial and temporal intelligence. Because the footage is mostly egocentric or selfie-style, models must construct a holistic understanding from partial views, which further stresses geospatial–temporal reasoning. The dataset comprises 200 travel vlogs filmed across Japan, a major inbound tourism destination, each accompanied by a manually annotated and double-reviewed visiting order graph.
@@ 2.1 Video Benchmarks
With the rapid advancement of MLLMs, recent video understanding benchmarks have increasingly prioritized evaluating models’ spatial and temporal reasoning capabilities. For instance, Ego4D grauman2022ego4dworld3000hours facilitates the evaluation of models’ comprehension of past and future events by utilizing curated temporal data, while HourVideo chandrasegaran2024hourvideo examines the performance of models in understanding extended-duration video content. VSI-Bench yang2025thinking assesses a model’s ability to infer 3D scene layouts from 2D video inputs, while OST-Bench lin2025ost evaluates spatial-temporal understanding by requiring models to explore and interpret information within a 3D space. CityGuessr kulkarni2024cityguessrcitylevelvideogeolocalization introduces a video-based benchmark for assessing geo-localization using driving videos, while UrbanVideo-Bench zhao2025urbanvideobenchbenchmarkingvisionlanguagemodels targets the embodied cognitive abilities of MLLMs within urban 3D environments using drone-collected footage. Nevertheless, most existing benchmarks primarily feature indoor scenarios or short-distance outdoor movements, lacking extensive long-distance traversal, such as inter-city journeys. Consequently, these benchmarks are insufficient for thoroughly evaluating the geospatial-temporal intelligence of MLLMs. In contrast, VIR-Bench specifically addresses this gap by comprehensively assessing video understanding capabilities across extended spatial (e.g., from Tokyo to Osaka, Hokkaido to Kyoto) and temporal (spanning multiple days) scales.
@@ 2.2 Itinerary Extraction
Researchers in natural language processing have substantially studied the task of extracting travel trajectories from text 10.1145/1869890.1869894; kaushik-etal-2017-making; ijgi10100710; yamamoto-etal-2025-graph. A representative study by yamamoto-etal-2025-graph introduces a visiting order graph designed to capture relationships among visited locations and provides a benchmark dataset for training and evaluating itinerary extraction models. In multimodal settings, PANG2011352 proposes a framework aimed at summarizing travelogues by integrating text and images from blogs. More recently, DBLP:conf/rectour/Rosa24 leverages MLLMs to perform structured entity extraction from travel videos, while zhuang2024vloggermakedreamvlog tackles the inverse problem by generating vlogs with diverse travel scenes. This study advances this line of work by providing, to our knowledge, the first systematic investigation into extracting and reconstructing itineraries directly from videos, establishing a new benchmark for video-centric geospatial and temporal understanding.
@@ 2.3 Itinerary Generation
The complexity of manual trip planning has driven research into automated itinerary generation. Initial approaches were often based on optimization problems like the Tourist Trip Design Problem and classic machine learning Gavalas2014; 36046891aa77462298f21dacee3279a6; 8731399; Carrillo2023SmartTurECO. More recently, Large Language Models (LLMs) have enabled more sophisticated and flexible frameworks chen2024travelagentaiassistantpersonalized; xie2024travelplanner. Current research trends include developing novel reasoning paradigms gui2025hypertree, creating hybrid systems that combine LLMs with classical planners delarosa2024trippaltravelplanningguarantees, and enhancing personalization through user model integration and interactive feedback singh-etal-2024-personal; chen2024travelagentaiassistantpersonalized; OTAKI2025126294.
To rigorously evaluate these methods, various benchmarks have been developed. Notable studies include real-world planning xie2024travelplanner, fine-grained spatio-temporal planning Chaudhuri2025TripCraft, and assessing personalization singh-etal-2024-personal.
@@ 3.1 Visiting Order Graph
We adopt and refine the definition of a visiting order graph introduced by yamamoto-etal-2025-graph. A simplified example is shown in Figure 2.
@@
To prevent cycles in the graph, we treat multiple visits to the same location as distinct nodes. Following yamamoto-etal-2025-graph, we also introduce a special “Overlap” edge to handle POIs that are geographically overlapping but cannot be represented through inclusion edges.
@@ 3.2 Data Annotation
After annotation, we retrieved detailed information of each POI including name, address and categories using Google Places API. Using the timestamped POIs, we then constructed a visiting order graph for each video (e.g., Figure 2). We also conducted a quality check of the generated graphs and corrected errors by rerunning the retrieval step manually when POIs were incorrectly annotated or matched. This pipeline yielded VIR-Bench, a dataset of 200 travel videos (100 in English and 100 in Japanese) paired with their corresponding visiting order graphs. In total, 3,689 POIs were identified across 43 of Japan’s 47 prefectures. Detailed annotation guidelines and dataset statistics are provided in Appendix A.
@@ 4.2 Benchmark Models
We evaluate the performance of mainstream MLLMs on VIR-Bench, including both open-weight models (VideoLLaMA3 zhang2025videollama3frontiermultimodal, LLaVA-Video zhang2024videoinstructiontuningsynthetic), InternVL3 zhu2025internvl3exploringadvancedtraining, Qwen2.5-VL bai2025qwen25vltechnicalreport) and proprietary models (GPT-4.1 openai2025_gpt41_systemcard, o4-mini openai2025_o3_o4_mini_systemcard, Gemini-2.5-Flash and Pro comanici2025gemini25pushingfrontier). All models are evaluated in a zero-shot setting. We use as many input frames as permitted by each model’s interface or pre-training setup; only the Gemini models accept audio input, and full details appear in Appendix B.1.
@@ 4.4 Main Results
When comparing models of different sizes within the same families (LLaVA-Video, InternVL3, Qwen2.5-VL), we observe steady, scale-driven gains on node prediction, whereas edge prediction shows a sharp jump; for example, transition F1 improves by about $16\times$ from Qwen2.5-VL-7B to Qwen2.5-VL-72B. This pattern reflects the task demands: node prediction is a localized, single-point task that primarily relies on geospatial knowledge encoded in the models, whereas edge prediction requires a holistic view of the itinerary and thus benefits more from larger models with stronger long-context and temporal reasoning. An exception is LLaVA-Video-72B, which shows minimal improvement over LLaVA-Video-7B in POI node prediction. This is likely due to the limited geographic coverage in the LLaVA-OneVision training data li2024llavaonevisioneasyvisualtask. In contrast, models like Qwen2.5-VL demonstrate strong geo-localization capabilities; in our internal evaluations, Qwen2.5-VL was able to accurately predict POI coordinates, indicating extensive pretraining on geographic data.
@@ 5.1 Task Definition
The agent system takes a list of POIs, a video and planning constraints as input. While this list of POIs could be the output of the node prediction task, we use the POI list from the video annotation in this experiment to isolate the evaluation from model performance. The video provides richer information for the planning process. The constraints consist of the number of companions, travel duration and travel budget inspired by previous work xie2024travelplanner.
@@ 5.3.2 Evaluation Setup
We also evaluate the system’s POIs selection. Using GPT-4o openai2024gpt4ocard, we extract the POIs mentioned in the generated plans and treat them as selected. We then compare selected versus unselected POIs in terms of their on-screen duration in the video and their Google Maps ratings.
@@ A.1.3 Trajectory Annotation
We use Vidat zhang2020vidat, a browser-based video annotation tool. Before starting, we strongly recommend watching the official 7-minute tutorial video.
@@ B.1 Model Cards
VideoLLaMA3-7B 180 - - zhang2025videollama3frontiermultimodal DAMO-NLP-SG/VideoLLaMA3-7B
LLaVA-Video-7B 64 - - zhang2024videoinstructiontuningsynthetic lmms-lab/LLaVA-Video-7B-Qwen2
@@
InternVL3-8B 64 - - zhu2025internvl3exploringadvancedtraining OpenGVLab/InternVL3-8B
@@
Qwen2.5-VL-7B 256 - - bai2025qwen25vltechnicalreport Qwen/Qwen2.5-VL-7B-Instruct
@@
GPT-4.1 256 - - openai2025_gpt41_systemcard gpt-4.1-2025-04-14
o4-mini 256 32768 - openai2025_o3_o4_mini_systemcard o4-mini-2025-04-16
Gemini-2.5-Flash 1 fps - 1 kbps comanici2025gemini_full gemini-2.5-flash